# Emergent Hawking Radiation and Quantum Sensing in a Quenched Chiral Spin Chain

### v1

@@ Abstract
We first consider the qubit strongly coupled to the chain. By tracing out the environment, we compute the exact reduced density matrix $\rho_{q}(t)$ (see Appendix C for details). In this regime, the spin chain acts as a finite many-body reservoir. The interaction induces energy exchange, causing the qubit populations to relax to a steady state on a timescale $\tau_{0}\propto(g_{1}^{2}+g_{2}^{2})^{-1/2}$ leading to rapid decoherence [^1] The decoherence function, defined as $D(t)=\prod_{p}|P_{ge}(p,t)|^{2}$, quantifies the magnitude of coherence present in the reduced density matrix.. As shown in Fig. 5, $D(t)$ decays sharply as the coupling between qubit and environment becomes stronger. However, the effective temperature $T_{ef}$ extracted from the steady-state populations does not coincide with the Hawking temperature $T_{H}$. Instead, $T_{ef}$ depends non-trivially on the coupling parameters $g_{1,2}$ and the global properties of the chain. In this strong-coupling limit, the qubit effectively thermalizes with the entire spin-chain environment rather than selectively probing the horizon-induced modes. Consequently, a strongly coupled qubit is not a faithful probe of HR.
@@
In this limit, the spin chain acts as a Markovian thermal bath. The reduced qubit dynamics are governed by rate equations determined solely by the bath spectral density $J(\omega)$ and the mode occupation Breuer and Petruccione [2007]:
@@
where the rates $\Gamma_{\uparrow,\downarrow}\propto J(\omega)(e^{\pm\omega/T_{H}}+1)^{-1}$ follow the Fermi-Dirac distribution. Unlike the strong coupling case, here the qubit is insensitive to microscopic details and probes only the thermal state of the Fermionic field. Solving the master equation confirms that the qubit relaxes to a stationary state $\rho_{ee}^{s}=(e^{\omega/T_{H}}+1)^{-1}$. Figure 6 illustrates the equilibration of the population ratio $\log(\rho_{ee}/\rho_{gg})$. The slope of this ratio in the steady state directly yields the Hawking temperature $T_{H}$. Thus, in the weak-coupling limit, the qubit functions as a faithful quantum thermometer, allowing for the experimental determination of the analogue Hawking temperature via simple population tomography.
@@
Conclusions and Outlook: In this work, we have established a robust analogue of HR in a chiral spin chain modeled as a sudden quantum quench. By deriving the response of both field-theoretic modes and quantum sensor probes, we have clarified the interplay between horizon formation, thermality, and the detectability.
@@
First, our analysis distinguishes between mathematical thermality and physical detectability. While idealized plane-wave modes recover a strictly thermal spectrum, we demonstrate that physically realistic probes — modeled as localized Gaussian wave packets — exhibit deviations from the Planckian distribution. These deviations arise from the finite localization of the detector, which imposes a frequency-dependent filtering analogous to greybody factors in gravitational systems. This result challenges the conventional view of HR as a structureless, information-free thermal bath; instead, it suggests that physical measurement processes naturally imprint non-thermal features onto the observed spectrum, potentially encoding information about the detector-horizon interplay at finite scales.
@@
Second, the universal nature of our results sheds light on the role of formation history. In our previous work, we showed that the geometric satisfaction of the Hoop conjecture is a necessary but insufficient condition for horizon formation in the percolation picture Jaiswal and Shankaranarayanan [2025], Bonnor [1983]. However, the present analysis reveals that the stationary HR is insensitive to these formation criteria. The emission statistics remain robustly Poissonian for both plane-wave and Gaussian probes. This insensitivity is not a limitation of the model but a manifestation of the universal kinematics of the near-horizon region. The radiation effectively washes out the memory of the formation scale (Hoop information), reinforcing the idea that HR acts as a late-time attractor that is independent of the microscopic details of the collapse.
@@
Third, we emphasize that our model operates strictly within the continuum limit of the spin chain. Consequently, it reproduces the continuous Hawking spectrum characteristic of semiclassical gravity but remains blind to discrete lattice effects such as the Bekenstein-Mukhanov spectrum Bekenstein and Mukhanov [1995]. The quantization of horizon area, which implies a discrete line spectrum for radiation, requires retaining the discrete lattice structure of the chain or use other probes like in fluid-gravity correspondence Cropp et al. [2017].
@@
Finally, we have identified the operational conditions required to measure this effect in a quantum simulation. We showed that a qubit acts as a faithful thermometer only in the weak-coupling regime, where it essentially probes the spectral density of the bath. In contrast, strong coupling induces complex many-body entanglements where the qubit thermalizes with the entire environment rather than the horizon specifically. This distinction is critical for future experimental realizations of analogue gravity in quantum spin systems, providing a clear protocol for distinguishing genuine Hawking temperatures from spurious environmental heating, especially in setups that mimic cosmological expansion Chandran and Fischer [2025].
@@
Acknowledgments: The authors thank Soumya Bhattacharya, Indranil Chakraborty, Mahesh Chandran, K. Hari, P. George Christopher, Tausif Parvez, Tapobrata Sarkar, and Archit Vidyarthi for discussions and comments on the earlier draft. The work is supported by SERB-CRG/2022/002348.
@@ Appendix A The Model and its Diagonalization
$H(t)=-\sum_{j=1}^{N}\frac{\mathcal{U}}{2}\left(\sigma_{j}^{x}\sigma_{j+1}^{x}+\sigma_{j}^{y}\sigma_{j+1}^{y}\right)+\Theta(t-t_{0})\sum_{j=1}^{N}\left[\frac{v}{4}\,\chi_{j}+(g_{1}\sigma_{q}^{z}+g_{2}\sigma_{q}^{x})\frac{\chi_{j}}{4}\right]~.$ (18)
Here, $\chi_{j}=\vec{\sigma}_{j}\cdot(\vec{\sigma}_{j+1}\times\vec{\sigma}_{j+2})$ denotes the three-spin chirality operator, $v$ is a dimensionless coupling constant, $\sigma_{q}^{x,z}$ are the Pauli matrices acting on the qubit, and $g_{1,2}$ represent the corresponding coupling strengths.
To diagonalize the Hamiltonian, we first observe that the operator $\mathcal{Q}\equiv g_{1}\sigma_{q}^{z}+g_{2}\sigma_{q}^{x}$ commutes with the total Hamiltonian, $\mathcal{Q},H(t)]=0$. As a result, $\mathcal{Q}$ is a conserved quantity and can be treated as a constant within each sector. Its eigenvalues are $\pm\sqrt{g_{1}^{2}+g_{2}^{2}}$, with corresponding eigenstates given by Quan et al. [2006], Liu et al. [2009]:
@@
$H(t)=\sum_{p}E_{\pm}(p)c^{\dagger}_{p}c_{p}~,\quad E_{\pm}(p)=-2\,\mathcal{U}\cos(p)+V_{e}^{\pm}\sin(2p)~,$ (21)
where $V_{e}^{\pm}=v\pm\sqrt{g_{1}^{2}+g_{2}^{2}}$. Here $p\in[-\pi,\pi)$ denotes the lattice quasi-momentum, $E_{\pm}(p)$ is the dispersion relation for the total Hamiltonian, and $c_{p}$ ($c^{\dagger}_{p}$) are fermionic annihilation (creation) operators obtained via Fourier transformation.
Appendix B Analog Hawking Radiation
@@ B.1 Plane Wave States
The covariant Dirac equation for massless spinor fields in curved spacetime is given by Jaiswal and Shankaranarayanan [2025]:
@@
$\tau=t-\int_{x_{0}}^{x}\frac{dy\,V_{e}^{\pm}}{\mathcal{U}^{2}(y)g(y)}~,\quad\mathrm{X}=\frac{i\sigma^{2}}{(V_{e}^{\pm})^{2}\,\mathcal{X}_{h}}\tan^{-1}\!\left(\frac{\mathcal{X}}{i\,\mathcal{X}_{h}}\right)~,\quad\mathcal{X}_{h}=\pm\sigma\sqrt{2(\sigma_{c}-\sigma)}~.$ (27)
@@
$\gamma_{c}^{\tau}=\frac{1}{g(\mathrm{X})}\begin{pmatrix}0&&1+\frac{V_{e}^{\pm}}{\mathcal{U}(\mathrm{X})}\\ 1-\frac{V_{e}^{\pm}}{\mathcal{U}(\mathrm{X})}&&0\end{pmatrix}~,\quad\gamma_{c}^{\mathrm{X}}=\frac{1}{g(\mathrm{X})}\begin{pmatrix}0&&-1-\frac{V_{e}^{\pm}}{\mathcal{U}(\mathrm{X})}\\ 1-\frac{V_{e}^{\pm}}{\mathcal{U}(\mathrm{X})}&&0\end{pmatrix}~$, (28)
@@
$W_{\tau}$ $=\frac{(V_{e}^{\pm})^{2}\,\mathcal{U}^{\prime}(\mathrm{X})}{2\,\mathcal{U}^{3}(\mathrm{X})g(\mathrm{X})}\gamma^{0}~,\quad W_{\mathrm{X}}=\frac{V_{e}^{\pm}\,\mathcal{U}^{\prime}(\mathrm{X})}{2\,\mathcal{U}^{2}(\mathrm{X})g(\mathrm{X})}\gamma^{0}~,$
$\mathcal{U}(\mathrm{X})$ $=\frac{1}{\sqrt{2\sigma}}\left[1+\left(1-\frac{1}{2\sigma{V_{e}^{\pm}}^{2}}\right)\cot^{2}\left(\frac{\mathrm{X}V_{e}^{\pm}\sqrt{2\sigma{V_{e}^{\pm}}^{2}-1}}{\sigma}\right)\right]^{-1/2}.$ (29)
The calculation of Hawking radiation crucially relies on tracing outgoing field modes backward in time toward the horizon. The above null coordinates play a central role in tracing outgoing modes backward toward the horizon and in evaluating the Bogoliubov coefficients responsible for Hawking radiation. As emphasized in Ford’s review Ford [1997], this backward propagation reveals the exponential redshift experienced by outgoing rays near the horizon:
@@
which in turn governs the nontrivial Bogoliubov mixing between positive and negative-frequency modes. The Bogoliubov coefficient $\beta_{ff^{\prime}}$ is defined by the Dirac inner product between the “in” and “out” modes is
$\beta_{ff^{\prime}}\equiv(\psi_{f}^{\text{out}},\,\psi_{f^{\prime}}^{\text{in}})_{D}=\int_{\Sigma}d\Sigma_{\mu}\bar{\psi}_{f}^{\text{out}}\gamma^{\mu}\psi_{f^{\prime}}^{\text{in}}~,$ (31)

### v2

@@ Abstract
[…]
where $P_{ee}(p,t)=1-P_{gg}(p,t)$ and the offdiagonal coherence term is $P_{ge}(p,t)=P_{eg}^{\ast}(p,t)$. The Rabi frequency is defined as $\Omega_{R}=2\sqrt{g_{1}^{2}+g_{2}^{2}}\sin{2p}$ and the total population is the product over all quasi-momenta $p$. As shown in Fig. 5, $D(t)$ decays sharply as the coupling between qubit and environment becomes stronger. However, the effective temperature $T_{ef}$ extracted from the steady-state populations does not coincide with the Hawking temperature $T_{H}$. Instead, $T_{ef}$ depends non-trivially on the coupling parameters $g_{1,2}$ and the global properties of the chain. In this strong-coupling limit, the qubit effectively thermalizes with the entire spin-chain environment rather than selectively probing the horizon-induced modes. Consequently, a strongly coupled qubit is not a faithful probe of HR.
@@
Here, the spin chain acts as a Markovian thermal bath. The reduced qubit density matrix are governed by master rate equations determined solely by the bath spectral density $J(\omega)$ and the mode occupation Breuer and Petruccione (2007):
@@
where the rates $\Gamma_{\uparrow,\downarrow}\propto J(\omega)(e^{\pm\omega/T_{H}}+1)^{-1}$ follow the Fermi-Dirac distribution. These rates yield the reduced qubit dynamics for the excited-state population $\rho_{ee}(t)=f_{D}(\omega)\left(1-e^{-J(\omega)t}\right),$ and the ground-state population is simply given by $\rho_{gg}(t)=1-\rho_{ee}(t)$, showing relaxation toward a stationary thermal state. Unlike the strong coupling case, here the qubit is insensitive to microscopic details and probes only the thermal state of the Fermionic field. Solving the master equation confirms that the qubit relaxes to a stationary state $\rho_{ee}^{s}=(e^{\omega/T_{H}}+1)^{-1}$. Figure 6 illustrates the equilibration of the population ratio $\log(\rho_{ee}/\rho_{gg})$. The slope of this ratio in the steady state directly yields the Hawking temperature $T_{H}$. Thus, in the weak-coupling limit, the qubit functions as a faithful quantum thermometer, allowing for the experimental determination of the analogue Hawking temperature via simple population tomography.
@@
Conclusions and Outlook: In this work, we have established a robust analogue of HR in a chiral spin chain modeled as a sudden quantum quench. Although the specific metric derived here arises from the spin chain couplings, the $\text{sech}(x)$ profile generates an effective geometry that shares the causal structure and multiple horizons of a Reissner-Nordström BH, making these insights applicable to astrophysical spacetimes. Furthermore, we observe that the HR is statistically insensitive to the ‘formation criteria’—the specific dynamical details of the quantum quench that created the horizon. While the spectral thermality fluctuates depending on whether the ‘in’ and ‘out’ vacua are probed via idealized plane waves or localized Gaussian wavepackets, the underlying emission remains robustly Poissonian. This memoryless statistical signature confirms that the core features of HR transcend both the microscopic UV completion and the specific dynamics of the gravitational collapse.
Our work uniquely bridges abstract quantum field theory in curved spacetime with modern quantum information science. By introducing a qubit as a global quantum sensor, we provide a concrete, operational framework for detecting horizon-induced correlations via collective modes. This clarifies which aspects of HR are genuinely geometric and which are tied to measurement dynamics, offering a novel protocol for probing macroscopic relativistic entanglement without requiring localized defect modeling.
@@
First, our analysis distinguishes between mathematical thermality and physical detectability. While idealized plane-wave modes recover a strictly thermal spectrum, we showed that physically realistic probes — modeled as Gaussian wave packets — exhibit deviations from the Planckian distribution. These deviations arise from the localization of the detector. This result challenges the conventional view of HR as a structureless, information-free thermal bath; instead, it suggests that measurement processes naturally imprint non-thermal features onto the observed spectrum, potentially encoding information about the detector-horizon interplay at finite scales.
@@
Second, the universal nature of our results sheds light on the role of formation history. In our previous work, we showed that the geometric satisfaction of the Hoop conjecture is a necessary but insufficient condition for horizon formation in the percolation picture Jaiswal and Shankaranarayanan (2025); Bonnor (1983). However, the present analysis reveals that the stationary HR is insensitive to these formation criteria. The emission statistics remain robustly Poissonian for both probes. This insensitivity is not a limitation of the model but a manifestation of the universal kinematics of the near-horizon region. The radiation effectively washes out the memory of the formation scale (Hoop information), reinforcing the idea that HR acts as a late-time attractor that is independent of the microscopic details of the collapse.
@@
Third, we emphasize that our model operates strictly in the continuum limit of the spin chain. Consequently, it reproduces the continuous Hawking spectrum characteristic of semiclassical gravity but remains blind to discrete lattice effects such as the Bekenstein-Mukhanov spectrum Bekenstein and Mukhanov (1995). The quantization of horizon area, requires retaining the discrete lattice structure of the chain or use other probes like in fluid-gravity correspondence Cropp et al. (2017).
@@
Finally, we have identified the operational conditions required to measure this effect in a quantum simulation. We showed that a qubit acts as a faithful thermometer only in the weak-coupling regime, where it essentially probes the spectral density of the bath. In contrast, strong coupling induces complex many-body entanglements where the qubit thermalizes with the entire environment rather than the horizon specifically. This distinction is critical for future experimental realizations of analogue gravity in quantum spin systems, providing a clear protocol for distinguishing genuine Hawking temperatures from spurious environmental heating, especially in setups that mimic cosmological expansion Chandran and Fischer (2025).
While the present work establishes the operational utility of a global quantum sensor, a natural future direction is to investigate a strictly localized probe coupled to a single lattice site $j_{0}$. Because such a localized coupling breaks translational invariance, this would require analyzing the horizon dynamics through the lens of open quantum systems impurity physics—such as a Caldeira-Leggett model with a localized defect to explore how exact spatial sampling affects the detected Hawking spectrum.
@@
Acknowledgments: The authors thank S. Bera, S. Bhattacharya, I. Chakraborty, K. Hari, P. George Christopher, T. Parvez, and T. Sarkar for discussions and comments on the earlier draft. We thank the anonymous referee for the constructive comments and suggestions, which helped improve the clarity and presentation of this manuscript. The work is supported by ANRF/ARG/2025/001514/PS.
@@ Appendix A The Model and its Diagonalization
$H(t)=-\sum_{j=1}^{N}\frac{u(j,j+1)}{2}\left(\sigma_{j}^{x}\sigma_{j+1}^{x}+\sigma_{j}^{y}\sigma_{j+1}^{y}\right)+\Theta(t-t_{0})\sum_{j=1}^{N}\left[\frac{v}{4}\,\chi_{j}+(g_{1}\sigma_{q}^{z}+g_{2}\sigma_{q}^{x})\frac{\chi_{j}}{4}\right]~.$ (19)
Here, $u(j,j+1)\equiv u(|(j+1)-j|)$ is a site-dependent coupling constant depending only on distance between the nearest-neighbor sites, $\chi_{j}=\vec{\sigma}_{j}\cdot(\vec{\sigma}_{j+1}\times\vec{\sigma}_{j+2})$ denotes the three-spin chirality operator, $v$ is a dimensionless coupling constant, $\sigma_{q}^{x,z}$ are the Pauli matrices acting on the qubit, and $g_{1,2}$ represent the corresponding coupling strengths. At $t=t_{0}$, a sudden quench leads to changing the XX spin-chain Hamiltonian to the post-quench Hamiltonian $H(t)$. Since the quench occurs on a time scale much shorter than the intrinsic dynamical scales of the system, the initial ground state does not match the ground state of the new Hamiltonian, leading to the generation of excitations and injected energy Quan et al. (2006); Jaiswal et al. (2022). This sudden quench is equivalent to a null shell collapse in the continuum limit of the above spin-chain model, triggering black hole formation in the emergent spacetime.
We first carry out the diagonalization of the environment Hamiltonian $H_{E}$ comprising of the standard $XX$ spin chain and the three-spin chirality term using Ref. Horner et al. (2023); Jaiswal and Shankaranarayanan (2025). We employ a Jordan–Wigner transformation, which provides an exact mapping between the spin-1/2 Pauli operators and a set of spinless fermionic operators defined as
$\sigma_{j}^{+}=\exp{\left(-i\pi\sum_{l=1}^{j-1}c_{l}^{\dagger}c_{l}\right)}c_{j}^{\dagger}~,\quad\sigma_{j}^{-}=\exp{\left(i\pi\sum_{l=1}^{j-1}c_{l}^{\dagger}c_{l}\right)}c_{j}~,\quad\sigma_{j}^{z}=1-2c_{j}^{\dagger}c_{j}~,$ (20)
where $\sigma_{j}^{\pm}=(\sigma_{j}^{x}\pm\sigma_{j}^{y})/2$, while $c_{j}$ and $c^{\dagger}_{j}$ are fermionic annihilation and creation operators, respectively satisfying anti-commutation relations $\{c_{j},c_{k}^{\dagger}\}=\delta_{jk}$ and $\{c_{j},c_{k}\}=\{c_{j}^{\dagger},c_{k}^{\dagger}\}=0$. Using this we can write the environment Hamiltonian
$H_{E}=\sum_{j=1}^{N}\left[-u(|(j+1)-j|)c^{\dagger}_{j}c_{j+1}+\frac{iv}{2}\left(c^{\dagger}_{j}c_{j+1}\sigma_{j+2}^{z}+c^{\dagger}_{j+1}c_{j+2}\sigma_{j}^{z}-c^{\dagger}_{j}c_{j+2}\right)\right]+H.C.~.$ (21)
[…]
$c_{j}=\frac{1}{\sqrt{N}}\sum_{p\in[-\pi,\pi)}e^{ipj}c_{p}~,\quad c_{j}^{\dagger}=\frac{1}{\sqrt{N}}\sum_{p\in[-\pi,\pi)}e^{-ipj}c_{p}^{\dagger}~,$ (22)
where the quasi-momentum $p=\frac{2\pi\lambda}{N}$ for $\lambda\in\mathbb{Z}$ yielding
$H_{E}=-\frac{1}{N}\sum_{p,\,p^{\prime}}\sum_{j=1}^{N}u(|(j+1)-j|)e^{i(p^{\prime}(j+1)-pj)}c_{p}^{\dagger}c_{p^{\prime}}+\sum_{p}\frac{iv}{2}e^{2ip}c_{p}^{\dagger}c_{p}+H.C.~.$ (23)
Rewriting the above Hamiltonian in the diagonalized form:
$H_{E}=\!\!\sum_{p\in[-\pi,\pi)}\!\!E(p)c^{\dagger}_{p}c_{p}~,\quad E(p)=-2\,\mathcal{U}\cos(p)+v\sin(2p)~,$ (24)
where $\mathcal{U}=u(|(j+1)-j|)$ and $E(p)$ is the dispersion relation of the environment Hamiltonian.
To diagonalize the total Hamiltonian $H(t)$ by incorporating the contribution from the qubit sector, we first observe that the operator $\mathcal{Q}\equiv g_{1}\sigma_{q}^{z}+g_{2}\sigma_{q}^{x}$ commutes with the total Hamiltonian, $\mathcal{Q},H(t)]=0$. As a result, $\mathcal{Q}$ is a conserved quantity and can be treated as a constant within each sector. Its eigenvalues are $\pm\sqrt{g_{1}^{2}+g_{2}^{2}}$, with corresponding eigenstates given by Quan et al. (2006); Liu et al. (2009):
@@
$H(t)=\sum_{p}E_{\pm}(p)c^{\dagger}_{p}c_{p}~,\quad E_{\pm}(p)=-2\,\mathcal{U}\cos(p)+V_{\cal E}^{\pm}\sin(2p)~,$ (27)
where $V_{\cal E}^{\pm}=v\pm\sqrt{g_{1}^{2}+g_{2}^{2}}$ and $E_{\pm}(p)$ is the dispersion relation for the total Hamiltonian.
The ground state energy density of the total Hamiltonian in the thermodynamic limit is:
$\rho_{\pm}=\frac{1}{2\pi}\int_{p:E_{\pm}(p)<0}dpE_{\pm}(p)\quad=\,\Biggl\{\begin{array}[]{@{}ll@{}}\frac{2\mathcal{U}}{\pi},&V_{\cal E}^{\pm}\leq\mathcal{U}\\ -\frac{\mathcal{U}^{2}+(V_{\cal E}^{\pm})^{2}}{V_{\cal E}^{\pm}\pi},&V_{\cal E}^{\pm}>\mathcal{U}\end{array}~.$ (28)
A discontinuity in the second-order derivative of $\rho_{\pm}$ with respect to the interaction strength parameter $V_{\cal E}^{\pm}$ signals a second-order QPT at $|V_{\cal E}^{\pm}|=|\mathcal{U}|$ Jaiswal and Shankaranarayanan (2025). For $|V_{\cal E}^{\pm}|>|\mathcal{U}|$, the system is in the chiral phase, while $|V_{\cal E}^{\pm}|<|\mathcal{U}|$ resembles the non-chiral phase.
Appendix B Analogue Hawking Radiation
@@ B.1 Plane Wave States
The covariant Dirac equation for massless spinor fields in curved spacetime is given by Jaiswal and Shankaranarayanan (2025):
@@
$\tau=t-\int_{x_{0}}^{x}\frac{dy\,V_{\cal E}^{\pm}}{\mathcal{U}^{2}(y)g(y)}~,\quad\mathrm{X}=\frac{i\sigma^{2}}{(V_{\cal E}^{\pm})^{2}\,\mathcal{X}_{h}}\tan^{-1}\!\left(\frac{\mathcal{X}}{i\,\mathcal{X}_{h}}\right)~,\quad\mathcal{X}_{h}=\pm\sigma\sqrt{2(\sigma_{c}-\sigma)}~.$ (34)
@@
$\gamma_{c}^{\tau}=\frac{1}{g(\mathrm{X})}\begin{pmatrix}0&&1+\frac{V_{\cal E}^{\pm}}{\mathcal{U}(\mathrm{X})}\\ 1-\frac{V_{\cal E}^{\pm}}{\mathcal{U}(\mathrm{X})}&&0\end{pmatrix}~,\quad\gamma_{c}^{\mathrm{X}}=\frac{1}{g(\mathrm{X})}\begin{pmatrix}0&&-1-\frac{V_{\cal E}^{\pm}}{\mathcal{U}(\mathrm{X})}\\ 1-\frac{V_{\cal E}^{\pm}}{\mathcal{U}(\mathrm{X})}&&0\end{pmatrix}~$, (35)
@@
$W_{\tau}$ $=\frac{(V_{\cal E}^{\pm})^{2}\,\mathcal{U}^{\prime}(\mathrm{X})}{2\,\mathcal{U}^{3}(\mathrm{X})g(\mathrm{X})}\gamma^{0}~,\quad W_{\mathrm{X}}=\frac{V_{\cal E}^{\pm}\,\mathcal{U}^{\prime}(\mathrm{X})}{2\,\mathcal{U}^{2}(\mathrm{X})g(\mathrm{X})}\gamma^{0}~,$
$\mathcal{U}(\mathrm{X})$ $=\frac{1}{\sqrt{2\sigma}}\left[1+\left(1-\frac{1}{2\sigma{V_{\cal E}^{\pm}}^{2}}\right)\cot^{2}\left(\frac{\mathrm{X}V_{\cal E}^{\pm}\sqrt{2\sigma{V_{\cal E}^{\pm}}^{2}-1}}{\sigma}\right)\right]^{-1/2}.$ (36)
The calculation of Hawking radiation crucially relies on tracing outgoing field modes backward in time toward the horizon. The above null coordinates play a central role in tracing outgoing modes backward toward the horizon and in evaluating the Bogoliubov coefficients responsible for Hawking radiation. As emphasized in Ford’s review Ford (1997), this backward propagation reveals the exponential redshift experienced by outgoing rays near the horizon:
@@
which in turn governs the nontrivial Bogoliubov mixing between positive and negative-frequency modes. In the context of Hawking radiation, Bogoliubov coefficients relate the vacuum defined in the distant past (“in” region) to that seen by observers in the far future (“out” region). They encode how gravitational collapse and horizon formation mix positive and negative-frequency modes, causing the initial vacuum to appear as a particle-filled state and leading to particle creation. These coefficients $\beta_{ff^{\prime}}$ and $\alpha_{ff^{\prime}}$ are defined by the Dirac inner product between the “in” and “out” modes
$\beta_{ff^{\prime}}$ $\equiv$ $(\psi_{f}^{\text{out}},\,\psi_{f^{\prime}}^{\text{in}})_{D}=\int_{\Sigma}d\Sigma_{\mu}\bar{\psi}_{f}^{\text{out}}\gamma^{\mu}\psi_{f^{\prime}}^{\text{in}}~,$
$\alpha_{ff^{\prime}}$ $\equiv$ $(\psi_{f}^{\text{out}},\,\psi_{f^{\prime}}^{{\text{in}}^{\dagger}})_{D}=\int_{\Sigma}d\Sigma_{\mu}\bar{\psi}_{f}^{\text{out}}\gamma^{\mu}\psi_{f^{\prime}}^{{\text{in}}^{\dagger}}~,$ (38)